[figure: 9.7 cm]

THE RESISTIVITY AND HALL EFFECT OF THE AMORPHOUS

Co_{x}Ni_{70−x}Fe_{5}Si_{15}B_{10} RIBBONS

K. Perzyńska^{a}, M. Biernacka^{a}, L. Dobrzyński^{a,b}, A. Malinowski, S. Tarasenko^{a} and P. Zaleski^{a}

aInstitute of Physics, University at Białystok, 15-424 Białystok, Lipowa 41, Poland balso Soltan Institute for Nuclear Studies, 05-400 Otwock-Świerk, Poland

The influence of structural disorder and magnetism on the resistivity of the amorphous Co_{x}Ni_{70−x}Fe_{5}Si_{15}B_{10} ribbons has been investigated. The resistivity vs. composition dependence shows parabolic-like behaviour. Low temperature logarithmic behaviour and resistivity minima are found. The temperatures of the minima vary as a function of the composition in similar way as the Curie temperatures do. The resistivity is proportional to T^{3/2} and T^{2} at temperatures T < 0.5 T_{c}, depending on the cobalt content. Both dependencies are considered to be of magnetic origin. Above T ≥ 0.5 T_{c} resistivity increases linear by with increasing temperature which is seemingly due to electronphonon scattering. Obviously, the chemical and positional disorder is responsible for the major part of conductivity.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : K. Perzyńska

Address for correspondence : Institute of Physics, University at Białystok 15-424 Białystok Lipowa 41, Poland

Email address : krystyna@alpha.uwb.edu.pl